# TIMeSynC: Temporal Intent Modelling with Synchronized Context Encodings for Financial Service Applications

### v2

@@ Abstract.
Users engage with financial services companies through multiple channels, often interacting with mobile applications, web platforms, call centers, and physical locations to service their accounts. The resulting interactions are recorded at heterogeneous temporal resolutions across these domains. This multi-channel data can be combined and encoded to create a comprehensive representation of the customer’s journey for accurate intent prediction. This demands sequential learning solutions. NMT transformers achieve state-of-the-art sequential representation learning by encoding context and decoding for the next best action to represent long-range dependencies. However, three major challenges exist while combining multi-domain sequences within an encode-decoder transformers architecture for intent prediction applications: a) aligning sequences with different sampling rates b) learning temporal dynamics across multi-variate, multi-domain sequences c) combining dynamic and static sequences. We propose an encoder-decoder transformer model to address these challenges for contextual and sequential intent prediction in financial servicing applications. Our experiments show significant improvement over the existing tabular method.
@@
journalyear: 2024conference: Eighteenth ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: The 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italyccs: Computing methodologies Machine learningccs: Information systems Recommender systemsccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Temporal data
@@ 3. Problem Statement
Figure 3. Illustrative Example: Data Flattening and Tokenization for the Encoder Context

### v3

@@ Abstract.
Users engage with financial services companies through multiple channels, often interacting with mobile applications, web platforms, call centers, and physical locations to service their accounts. The resulting interactions are recorded at heterogeneous temporal resolutions across these domains. This multi-channel data can be combined and encoded to create a comprehensive representation of the customer’s journey for accurate intent prediction. This demands sequential learning solutions. NMT transformers achieve state-of-the-art sequential representation learning by encoding context and decoding for the next best action to represent long-range dependencies. However, three major challenges exist while combining multi-domain sequences within an encoder-decoder transformers architecture for intent prediction applications: a) aligning sequences with different sampling rates b) learning temporal dynamics across multi-variate, multi-domain sequences c) combining dynamic and static sequences. We propose an encoder-decoder transformer model to address these challenges for contextual and sequential intent prediction in financial servicing applications. Our experiments show significant improvement over the existing tabular method.
@@
copyright: acmlicensedjournalyear: 2024conference: Eighteenth ACM Conference on Recommender Systems; Oct 14–18, 2024; Bari, Italyccs: Computing methodologies Machine learningccs: Information systems Recommender systemsccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Temporal data
@@ 3. Problem Statement
Figure 3. Illustrative Example. Data Flattening and Tokenization of the Encoder Context